OpenStudy (anonymous):

Factor fully: (x^2-y-^2)^2 ... Perfect squares or difference of squares?? Where do I go from here?? :/

hartnn (hartnn):

perfect square formula will not give you 'factors' (a+b)^2 = a^2 +2ab +b^2 no factors difference of squares will give you factors (A^2-B^2) = (A+B)(A-B) A+b and A-b are two factors :)

hartnn (hartnn):

so first try to factor x^2 - y^-2 using A^2-B^2 can you ?

OpenStudy (ankitshaw):

x^2 - y^2 is same as (x-y)(x+y) so after this you can proceed

OpenStudy (ankitshaw):

\[(x^2 -y^2)^2 = (x+y)^2(x-y)^2\]